### v1

# Towards Terminology Management Automation for Arabic

Abstract
This paper presents a method and supporting tools for automation of terminology management for Arabic. The tools extract lists of parallel terminology matching terms in foreign languages to their Arabic counterparts from field specific texts. This has significant implications as it can be used to improve consistent translation and use of terms in specialized Arabic academic books, and provides automated aid for enhancing cross-lingual text processing. This automation of terminology management aims to reduce processing time, and ensure use of consistent and correct terminology.
The extraction takes advantage of naturally occurring term translations. It considers several candidate phrases of varying lengths that co-occur next to the foreign terms.
Then it computes several similarity metrics, including lexicographic, phonetic, morphological, and semantic ones to decide the problem. We experiment with heuristic, machine learning, and ML with post processing approaches. This paper reports on a novel curated dataset for the task, an existing expert reviewed industry parallel corpora, and on the performance of the three approaches. The best approach achieved $94.9\%$ precision and $92.4\%$ recall.
Towards Terminology Management Automation for Arabic
Mahdi Nasser  and Laura Sayyah  and Fadi A. Zaraket Arab Center for Research and Policy Studies, Doha {mnasser,lsayyah,fzaraket} @dohainstitute.edu.qa
1 Introduction
Translation of academic literature is essential to knowledge transfer. Linguistic and structural complexities exist. These affect consistency and clarity of translated material and may hinder knowledge accessibility. Terminology management concerns the complexities involved in (i) specifying the term that should be used upon the introduction of a new concept to a language, and (ii) and enforcing the use, hence (iii) ensuring consistency and correctness.
The first and most important step is to build TermBase, a large database of parallel terms. Authors can utilize TermBase to enhance their writing processes when faced with a new concept; be it a foreign term to be translated to Arabic.
Recent advances in natural language processing and computational linguistics have paved the way for automated systems that assist in cross-lingual text processing. However, existing systems, including LLMs, often struggle with the nuances of terminology in cross-lingual specific expert fields, where precision is paramount. This paper introduces a novel methodology and a supporting framework to extract terminology mappings from existing Arabic books, a rich source of terms when building a TermBase.
When a foreign term occurs in Arabic books, usually between parentheses, it is preceded by its Arabic translation. Two problems arise. The exact scope of the Arabic term is ambiguous, and often times authors (even the same author) use Arabic terms inconsistently for the same non-Arabic ones. Editors and reviewers, who are usually expert researchers, are required to manually solve these laborious and time consuming tasks during lengthy editing iterations.
Our work helps automate these two tasks to reduce review and edit efforts and help reviewers and editors concentrate on more important research and knowledge generation tasks.
The problem is formulated as follows: given a sequence $s=\langle(f)w_{1}w_{2}w_{3}\ldots w_{n}\rangle$ representing a sentence with a foreign term $f$, denote candidate terms as the set of strings $s_{i}=w_{1}\ldots w_{i}$ ending before term $f$, $1\leq i\leq n$ as a candidate term. The goal is to identify the target candidate term $s_{t}$ best matching $f$. Our tool extracts linguistic features for each candidate match $(s_{i},f)$. The tool interpolates the features and aggregates their values into a resulting score that ranks the candidate terms.
By addressing the challenges of terminology mapping, this research moves us closer to a more integrated community where language barriers are significantly reduced, facilitating a greater exchange of knowledge and ideas across languages and cultures.
We make verified parts of the constructed dataset available online for the research community. To the best of our knowledge, this approach to TermBase extraction for Arabic is novel, and no comparable datasets currently exist.
1.1 Motivating example
Consider the following two contexts. They contain the same underlined french term l’ethnocentrisme with different translations.
  • Context 1:
    - (l’ethnocentrisme) , : ¡(…) نّ ¿.
    Gloss: "His sharp criticism of ethnocentrism (l’ethnocentrisme) does not turn into cultural relativism, but rather into rational universality."
  • Context 2:
    اً (l’ethnocentrisme) ¡ ¿, ذّ ¡ ¿.
    Gloss: "He strongly criticized the ethnocentric tendency (l’ethnocentrisme) that seeks to replace primitive human ideas about the world and society with our European ideas."
Candidates for Context 1:
  • - Gloss: "And his sharp criticism of ethnocentrism does not turn into…"
  • - Gloss: "His sharp criticism of ethnocentrism turns into…"
  • - Gloss: "His sharp criticism of ethnocentrism"
  • - Gloss: "The sharp ethnocentrism"
  • -    (target term) Gloss: "Ethnocentrism"
  • Gloss: "Ethnicity"
Candidates for Context 2:
  • اً Gloss: "He strongly criticized the ethnocentric tendency"
  • اً Gloss: "The strong ethnocentric tendency"
  • Gloss: "The ethnocentric tendency"
  • (target term) Gloss: "Ethnocentrism"
  • Gloss: "Centralism"
The two contexts and the extracted translation candidate terms showcase the inconsistency across contexts, and also expose potential adoption of different candidates by readers of each context. Thus amplifying the problem.
This paper makes the following contributions:
  • We propose a novel method for terminology mapping that leverages naturally occurring term translations, combining NLP and data driven techniques.
  • We create a unique annotated dataset for terminology extraction, which can serve as a valuable resource for the research community.
  • We evaluate and compare three approaches: heuristic, machine learning, and machine learning with post-processing demonstrating the effectiveness of these methods in addressing terminology mapping challenges. The results are impressive and convincing for deployment. The final model runs locally, requires no LLM interaction, and thus requires taking no IP risks from Arabic publishing houses.
We discuss TermBase creation in Section 2, the methodology in Section 3, briefly cover related work in  4, present and discuss our results in Section 5. We then discuss future directions and conclude.
2 Dataset creation
The source dataset is derived specifically from books published by our publishing house. These books span multiple academic fields, offering a diverse range of terminologies. The dataset includes terms in foreign languages (usually enclosed in parentheses) alongside their surrounding Arabic text.
The labeling process involved extracting candidate Arabic terms preceding each foreign term and annotating them with a binary label: True if the candidate is the correct translation of the foreign term, and False otherwise. To streamline this process, a draft of labels was generated using heuristic-based rules, and were subsequently reviewed to correct the errors.
2.1 Source data selection
To ensure diversity and relevance, we selected texts from multiple domains. The candidate terms were extracted by identifying all strings of words, with a certain relative length with respect to the foreign term, immediately preceding a parenthetical foreign term. This ensures that potential translations are almost always captured for further analysis.
2.2 Annotated dataset construction
Each extracted candidate term, along with its associated foreign term, was presented as an individual instance. The features of this instance, extracted using NLP tools, are described in section 3.
Subsequently, a draft of the labels was generated using the heuristics detailed in subsection 5.1 to facilitate the labeling process. This draft was then reviewed to ensure the accuracy of the labels.
3 Methodology
[…]
Figure 1: An overview of the process
Figure 1 illustrates the flow of our methodology. First, the tool extracts the candidates from a context (usually a paragraph) with a foreign term. Then, the feature vector of each candidate is computed. Finally, each vector is fed to a model which outputs a score that helps in choosing the target translation of the source foreign term.
3.1 Feature extraction
Given a candidate $s_{i}$, its foreign term $f$, and its context $c$, we extract features of the candidate that are going to be used later to predict which candidate is the target term. These features with their possible values are listed in Table 1 and explained in the following subsections.
Table 1: Features extracted for each candidate
Feature Value
Semantic similarity [0, 1]
Translation lexical
similarity [0, 1]
Transliteration lexical
similarity [0, 1]
Entity {PER, LOC, ORG, MISC}
POS {adj, adv, conj, misc, noun, noun_prop, part, prep, pron, verb}
Phonetic similarity {False, True}
3.1.1 Semantic similarity
We utilized the sentence transformer LaBSE Feng et al. (2022) to compute the embeddings of the source term and candidate term.
After that, we compute semantic similarity as the cosine similarity between these 2 vectors Salton et al. (1975). This might seem sufficient but it’s not, which is due to the high similarity between candidates (given any two candidates, one of them is a substring of the other). Therefore, we need more features to handle peculiar cases.
Table 2: Semantic similarity example
Candidate Semantic similarity
Gloss: "Anne Herzberg" 0.635
Gloss: "Herzberg" 0.66
Table 2 shows how the semantic similarity alone score favors the $2^{nd}$ candidate, which missed the first name in the foreign term and is incorrect according to experts.
3.1.2 Translation lexical similarity
Translation lexical similarity uses automated translation of the foreign term and measures a lexical distance between that and the candidate terms. The distance of choice is the Levenshtein Levenshtein (1966) distance which computes how many changes are needed for the terms to be identical. We use a ratio with respect to length to normalize the distance across candidates.
Example: suppose the foreign term is “London School of Economics and Political Science". Automatic translation returns
$t=$
Gloss: "London School of Economics and Political Science."
We have the candidate:
$s=$
Gloss: "London School of Economics and Political Science."
The Levenshtein distance between $s$ and $t$ is 15. The ratio is calculated as
$\frac{\text{len}(t)+\text{len}(s)-15}{\text{len}(t)+\text{len}(s)}=0.79$
3.1.3 Transliteration lexical similarity
This feature captures the cases where the target Arabic term is just a transliteration of the source term by computing the Levenshtein ratio between the transliteration of the source term and the candidate. This is helpful in the case of translating proper nouns that should supposedly map phonetically.
Example: suppose the foreign term is "Ehud Prawer". After transliterating it, we get
$t=$
Gloss: "Ehud Brower"
We also have the candidate with a hamza instead of an alef
$s=$
Gloss: "Ehud Brower"
The Levenshtein distance between $s$ and $t$ is 1. The ratio is calculated as
$\frac{\text{len}(t)+\text{len}(s)-1}{\text{len}(t)+\text{len}(s)}=0.95$
3.1.4 Entity
A named entity is a real-world object or concept that can be clearly identified and categorized with a proper name.
Candidates that are detected as named entities have a higher probability of being a term. We use an open source Arabic NER model  Jarrar et al. (2022) to detect entities in Arabic context. Also, NER is used on the source term to favor candidates having the same entity type; we used the spaCy  Honnibal et al. (2020) model xx_ent_wiki_sm in particular.
3.1.5 Part of speech
A part of speech (POS) is a category of words with similar grammatical properties. Common parts of speech include nouns, verbs, adjectives, adverbs, pronouns, prepositions, conjunctions, and interjections. Terms are mostly noun phrases and rarely start with verbs.
We used POS tagging tools Obeid et al. (2020) with its maximum likelihood expectation morphological disambiguator to detect the POS of the first word occurring in a candidate.
3.1.6 Phonetic similarity
As illustrated in Table 3, we use the soundex algorithm Russell and Odell (1918) to determine whether the source and candidate terms sound the same. This serves a similar purpose to transliteration lexical similarity; however, practicaly characterized with higher precision and lower recall.
Table 3: Phonetic similarity examples
Source term Candidate Phonetically similar
Regavim Gloss: "Regavim" True
Lockheed Martin Gloss: "Lockheed Regavim" True
Lockheed Martin Gloss: "like Lockheed Regavim" False
4 Related work
Besides bilingual dictionaries, efforts have been deployed to develop lexical sources from several original languages into Arabic. Arabic WordNet Global WordNet Association for instance relies on several open and restricted domain sources to develop a lexical English-Arabic database in different fields. Other more specialized glossaries have been developed such as the work of Hossam Mahdi’s Glossary of Terms for the Conservation of Cultural Heritage in Arabic Alphabetical Order Mahdy . Furthermore, glossaries offering support to interpreters Sharif have also been established.
The Arabic digital ontology Jarrar (2022) includes digitized Arabic dictionaries as denoted by dictionary authors. It requires API access for term lookup.
However, these are not complete, need to be updated manually regularly for novel terms, and are not easy to integrate in the editing process.
5 Results
Table 4 shows statistics of the Arabic books we have processed and the annotated dataset we have built to train and evaluate our work.
Table 4: Statistics from processed books
Arabic books 496
Unique foreign terms extracted 58,300
All occurrences of foreign terms 83,516
Total candidates 301,326
Average number of candidates per occurrence 3.61
Number of annotated books 9
Number of foreign terms in annotated dataset 3,595
Number of candidates in annotated dataset 15,865
5.1 Heuristics
In this approach, we came up with the following formula after multiple iterations of observing the results and fine-tuning:
$s=S_{L}+S_{S}+S_{E}+S_{P}+S_{POS}$
1. Lexical score $S_{L}$:
$S_{L}=\mathbb{1}\left(l_{1}\geq\tau\right)\cdot(1.2\cdot l_{1})+\mathbb{1}\left(l_{2}\geq\tau\right)\cdot l_{2}$
where $l_{1}$ is the lexical similarity with the translation, $l_{2}$ is with the transliteration, and $\tau=0.7$ is the similarity threshold.
2. Semantic score $S_{S}$:
$S_{S}=1.45\cdot S$
where $S$ is the semantic similarity score.
3. Entity score $S_{E}$:
$S_{E}=\begin{cases}0.7,&e\neq\text{"none"}\text{ and }e=e_{s}\\ 0.3,&e\neq\text{"none"}\text{ and }e\neq e_{s}\\ 0,&\text{otherwise}\end{cases}$
where $e$ is the candidate’s entity label and $e_{s}$ is the source term’s entity label.
4. Phonetic score $S_{P}$:
$S_{P}=\begin{cases}1,&\text{if phonetically similar}\\ 0,&\text{otherwise}\end{cases}$
5. Part of speech score $S_{POS}$:
$S_{POS}=\begin{cases}-1,&\text{if pos}\in\{\text{"verb"},\text{"prep"},\text{"conj"}\}\\ 1,&\text{if pos}\in\{\text{"noun"},\text{"noun\_prop"}\}\\ 0,&\text{otherwise}\end{cases}$
$\mathbb{1}(A)$ is the indicator function:
$\mathbb{1}(A)=\begin{cases}1,&\text{if }A\text{ is true}\\ 0,&\text{otherwise}\end{cases}$
After that, we select the candidates (of the same occurrence of a foreign term) with the highest scores within a maximum deviation (taken 0.03 here). Table 5 shows the results achieved by the heuristics approach on the whole dataset. Note that fine-tuning was done only by observing the training data. These decent results provided a good starting point that later helped in the labeling process of the data.
5.2 Machine learning
An instance of the dataset is the features of a candidate, and a label that is True if the candidate is the target candidate and False otherwise.
We also augmented the feature space with the following: semantic similarity rank within the candidates related to the same occurrence of a foreign term, semantic similarity difference with the candidate ranked $1^{st}$, translation lexical similarity rank, translation lexical similarity difference, transliteration lexical similarity rank, transliteration lexical similarity difference. The intuition behind the augmentation is to provide comparative context to the ML model with respect to other candidates.
Let $x\in$semantic similarity, translation lexical similarity, transliteration lexical similarity},
$A_{i}^{x}$ = Array of values of feature $x$ of candidates related to the same occurrence $i$ of a term, sorted by decreasing order.
For each candidate j:
[…]
where the candidate j belongs to the occurrence occ(j).
As shown in Table 4, We annotated 15,865 candidates to build a training dataset. Then we trained a binary classifier using the Auto-WEKA package on the Weka Workbench Frank et al. (2016). We allowed Weka to run for several hours on 8 threads, which resulted in the following:
{strip}
Best classifier: weka.classifiers.trees.RandomForest
Arguments: [-I, 131, -K, 28, -depth, 0]
Attribute search: weka.attributeSelection.GreedyStepwise
Attribute search arguments: [-C, -R]
Attribute evaluation: weka.attributeSelection.CfsSubsetEval
Attribute evaluation arguments: [-M]
Metric: errorRate
Table 5 shows the results.
We also tried ranking candidates by partitioning the testing dataset into groups of candidates related to the same occurrence of a term, and ranking them based on the value of their prediction instead of fixing a threshold, and then the candidate with the highest rank was chosen as the target term, and the others where detected as False. This achieved a 92% precision and 98.4% recall on the total dataset, which is worse than the 99% precision and 98.5% recall achieved by the direct application of the Random Forest on the total dataset.
Table 5: Comparison of Heuristics and ML Results
Category Metric Training Testing
Heuristics Precision 87%
Recall 89%
F1 Score 88%
ML Precision 100% 94.9%
Recall 100% 92.4%
F1 Score 100% 93.7%
Dataset Size 12,692 (80%) 3,173 (20%)
5.3 Discussion and error analysis
Upon further inspection, it was observed that most of the classification errors made by the ML model involved cases that required expert judgment to discern their correctness. This highlights an advantage of the model: its errors occur primarily on inherently difficult cases, which are challenging even for experts. Table 6 shows an example.
Table 6: Selecting a wrong candidate for the foreign term SA (Gloss: "Sturmabteilung")
Candidate Prediction Label
Gloss: "Storm Battalion" False True
Another cause of errors are peculiar cases where there is one letter difference between the correct translation and the selected candidate, which is caused by Arabic prepositions linked to a word. This is usually handled through pre-processing or providing both candidates to the model for it to decide on one, which might fail in some cases. Table 7 shows an example.
Table 7: Selecting a wrong candidate for the foreign term Verteilungsstellen (Gloss: "distribution points")
Candidate Prediction Label
Gloss: "distribution entities" True True
Gloss: "by distribution entities" True False
6 Conclusion
This paper introduced methods for extracting terminology mappings between Arabic and foreign languages from Arabic books. The approach achieves useful results, contributing to automated Terminology Management and text processing. The created dataset and existing corpora further enrich the research community, advancing cross-lingual knowledge exchange.
Limitations
Some terms have subtle translations that are inherently complex and challenging to identify, even for human experts. Therefore, the tool is designed to assist rather than replace expert judgment, effectively handling simpler terms automatically while supporting experts with more difficult cases.
Ethics statement
The data was collected and used with the appropriate approvals of the intellectual property owners. All results are reported following best academic standards and practices.